OpenStudy (anonymous):

z varies inversely as w z=10 when w=0.3 find z when w=40 please show all your work... so i can properly understand how you got the answer

OpenStudy (anonymous):

Set up the equation first: z = k/w Next plug in what is given to find k: z = k/w 10 = k/.3 Multiply both side by .3 10*.3 = k 3 = k Now rewrite the equation with the k we found: z = 3/w Last step; find z when w = 40 z = 3/40 = 0.075